OpenStudy (cameronlovesmorgan):

One painter can paint a house in 12 hours. The other painter can paint a house in 8 hours. a. Write an equation in terms of time, t, for the amount of time it would take both painters to paint the house together.

OpenStudy (eliesaab):

\[\frac 1 {12}+ \frac 1 8\] is what fraction of the house will be painted in one hour

OpenStudy (eliesaab):

\[ \frac 1 t \] is what fraction of the house will be painted in one hour

OpenStudy (eliesaab):

What is the equation?

OpenStudy (cameronlovesmorgan):

1 t --- + ---? 12 8

OpenStudy (eliesaab):

No\[ \frac 1{12} + \frac 1 8=\frac 1 t \]

OpenStudy (eliesaab):

For problems like these, visit http://www.saab.org/mathdrills/workcompletion.cg

OpenStudy (cameronlovesmorgan):

how would I solve?

OpenStudy (eliesaab):

\[ \frac{1}{12}+\frac{1}{8}=\frac{5}{24}=\frac 1 t \]

OpenStudy (eliesaab):

\[ t= \frac {24}5= 4 \frac 4 5= 4 \ hours \ and \ 48 \ min\]

OpenStudy (eliesaab):

Again practice on http://www.saab.org/mathdrills/workcompletion.cgi